OpenStudy (anonymous):

sqrt(x)+sqrt(a+x)=(2a/sqrt(a+x))??

OpenStudy (anonymous):

seems unlikely

OpenStudy (anonymous):

My text gives the answer x=3/a

OpenStudy (anonymous):

?? what exactly is the question?

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{a+x}=\frac{2a}{\sqrt{a+x}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@satellite73